Identify the following conic and hence reduce it in its standard form:

5x2-24xy-5y2+4x+58y-59=0


1
Expert's answer
2021-07-30T04:05:01-0400

"5x^2-24xy-5y^2+4x+58y-59=0"

Compare with

"ax^2+bxy+cy^2+dx+ey+f=0\\
a=5; b = -24; c=-5; d=4; e= 58; f=-59\\
Now \\
b^2-4ac\\
(-24)^2-4*5*(-5)= 676\\
b^2-4ac > 0 \\
Hence \space \space a \space hyperbola \\
The \space equation \space is \\
\frac{x^2}{a^2}-\frac{y^2}{b^2}=1\\
\frac{x^2}{5^2}-\frac{y^2}{(-24)^2}=1\\
\frac{x^2}{25}-\frac{y^2}{576}=1\\"